OpenStudy (anonymous):

reduction of order find second solution y2 y”+9y=0 y1=sIn3x

OpenStudy (goformit100):

@Wislar

OpenStudy (anonymous):

I believe what they want you to do is something of the form: \[y_1=\sin(3x)\] where y_1 denotes a solution, then by the principle of superposition: \[y=c \sin(3x)\] will also be a solution. Now to perform reduction of order, let c be a rather a function of x than a constant, such that: \[y_2=c(x) \sin(3x)\]